Course Contents» Set 02 (01/29 Tu 10 PM) Earths rotation TimerNotesvaluate Feedback Print It takes 23 hours 56 minutes and 4 seconds for the earth to make one revolution (mean sidereal day). What is the angular speed of the earth? 7.29x10 rad/s You are correct. Your receipt no. is 155-6866:① Previous Tries Assume the earth is spherical. Relative to someone on the rotation axis, what is the linear speed of an object on the surface if the radius vector from thé center of the earth to the object makes an angle of 42.0° with the axis of rotation. The radius of the earth is 6.37x103 km 3.11x102 m/s You are correct. Your receipt no. is 155-6424 Previous lries What is the acceleration of the object on the surface of the earth in the previous problem? What is the acceleration of an object undergoing uniform circular motion? Submit Answer Incorrect. Tries 10/12 Previous Tries Post Discussion Send Feedback
Answer #1

Solution) This problem is based on centripetal force

F = mv^2 /R

But F=ma

ma = mv^2 /R

a = v^2 /R   

Here v is linear speed ,v = 3.11×10^2 m/s from above

Radius = 6.37×10^(3)km = 6.37 ×10^(6) m ( 1km=1000m)

a = (3.11×10^2)^2 / ( 6.37×10^6)

a = 0.015 m/s^2